40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 133 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  10384   Thu Aug 14 15:10:47 2014 AndresUpdateIMCCalculation for the input mode cleaner

Quote:

Can you please give us some more details on how this design was decided upon? What were the design considerations?

It would be nice to have a shorter path length for WFS2. What is the desired spot size on the WFS? How sensitive are they going to be to IMC input alignment? Are we still going to be recentering the WFS all the time?

 I did the calculation, and I reduced the beam Path. In my calculation, I restricted the waist size at the WFSs to be between 1mm-2mm also the other parameter is that the Gouy Phase different between the WFSs have to be 90 degrees. I also try to minimize the amount of mirrors used. I found the Gouy phase to be 89.0622 degrees between the WFSs and the following table shows the solution that I got from a la mode:

 

  label                         z (m)                   type               parameters         
    -----                         -----                    ----                  ----------         
    MC1                    0                        flat mirror           none:            
    MC3                    0.1753               flat mirror           none:            
    MC2                   13.4587              curved mirror    ROC: 17.8700 (m)       
    Lens1                 28.8172              lens                   focalLength: 1.7183(m)
    BS2                    29.9475              flat mirror           none:            
    First Mirror         30.0237              flat mirror           none:            
    Lens3                 30.1253              lens                  focalLength: -0.100 (m)
    Lens2                 30.1635              lens                 focalLength: 0.1250(m)
    WFS1                 30.2269              flat mirror         none:            
    Second Mirror    30.2650              flat mirror         none:            
    Third Mirror       30.5698              flat mirror         none:            
    Lens4                30.8113              lens                  focalLength: -0.075 (m)
    WFS2                31.0778              flat mirror         none:     
       

In the first image attached below is the a la mode solution that show the waist size in the first WFS, and I used that solution to calculate the solution of the waist size for the second WFS, which is shown in figure 2. I photoshop a picture to illustrate how the new setup it supposed to look like. 

Attachment 1: SolutionForTheModeCleanerSetup00.png
SolutionForTheModeCleanerSetup00.png
Attachment 2: SolutionForTheModeCleanerSetup11.png
SolutionForTheModeCleanerSetup11.png
Attachment 3: PossibleSetupForModeCleaner.PNG
PossibleSetupForModeCleaner.PNG
Attachment 4: alaModeSolution.zip
  10383   Thu Aug 14 14:58:03 2014 JenneUpdateGeneralUpdated game plan

2014_Aug_14.pdf

(Updated as of 4pm)

  10382   Thu Aug 14 02:51:46 2014 JenneUpdateGeneralGame plan

[Jenne, Rana]

* Decided that earlier mode hop scan won't give us the information that we were hoping for.  We need to think about where we can actually see the frequency change.  Can we use the IR beatnote that we will soon have to do this?  We'd only be able to scan one laser temp at a time, but that's okay.  Leave, say, the PSL temperature alone, and scan one of the end laser temps.  Using the PSL as the reference, we will be able to see if the frequency of the end laser goes crazy and jumpy as we pass through a certain temp.  Then, repeat while holding the end laser constant and scan the PSL.  Thoughts?

* Meditated on PSL oplev servo, but I need to make a Matlab script that can evaluate different loops according to a cost function based on elog 9690

* Aligned IFO to IR, then greens to arms (got back to 0.9 for GTRY, but only about 0.5 for GTRX, with the PSL green shutter closed).  Then aligned green beams on the PSL table, since the PSL green pointing had changed a bit from Q's crystal alignment tweak-up earlier today.  Beatnotes are nice and big (see elog 10381 - The Yarm is the larger beatnote, and the Xarm is the smaller one.)

* Was not able to lock ALS comm/diff and hold long enough to get both arms to IR resonance.  Also, saw that TRY's RIN was more than 50%(!!!).  We took a look, and there seems to be much more low frequency noise than there was when the spectrum in the control room was taken for the multicolor metrology paper:

Y_ALS_FINE_PHASE_OUT_noisy_13Aug2014.pdf

* Tried to balance the ALS comm/diff input matrix, with not a lot of success.  First of all, it looks like the Xarm has overall about 10 times more noise!  We were exciting MC2 in position (~88 Hz, about 130 counts I think), and then looking at DARM_IN1 for the peak.  When DARM_IN1 was just one of the 2 ALS error signals (i.e. one matrix element set to zero), versus when both matrix elements were set to 1, we saw a factor of only about 3 in reduction of the peak height.  We were hoping to have better cancellation of this pure CARM signal in the DARM channel.  The Xarm green PDH loses lock every ~5 or 10 minutes, and when we relock it, this cancellation seems different, so we want to try again tomorrow when the ALS is locked on comm / diff, rather than just the free running ALS that we have now.  Although, if the balance of the input matrix changes lock-to-lock, we may need to consider redoing the green PSL table layout so we get a pure DARM beatnote signal like they have at the sites.

* We want to change how the watch script for ALS works, although this is a low-priority task.  Rather than looking at the control signal, we should maybe look at the sum of all the coil outputs, multiplied by a pendulum TF, and use that as a rough displacement sensor.  We want to be careful of pushing too hard at low frequencies, but we want to allow higher frequency actuation without having the watch script shut things down.

* Also, I should put on the to-do list the revamp of the ALS find IR resonance script.

  10381   Wed Aug 13 23:58:49 2014 ranaHowToComputer Scripts / ProgramsHP8591E spectrum analyzer remote scan

Quote:

The script for running continuous scans on HP 8591E spectrum analyzer is located at scripts/general/netgpibdata/HP8591E_contdScan.py

There was no such script in the directory when I looked today, but I found one called HP8591E. Of course, it didn't run because it hadn't been tested from the scripts directory and pointed to some /users/nichin/ stuff.

I modified a couple of lines and then committed it and the default .YML parameter file to the SVN. It runs and produces plots continuously from the scripts directory.

*** also, as you can see, we have mostly recovered the green beat amplitudes after yesterday's FLL attack on the ALS ***

Attachment 1: HP8591E_View.pdf
HP8591E_View.pdf
  10380   Wed Aug 13 23:08:17 2014 ranaUpdateIOOFSS box TFs

As EQ pointed out recently, we're injecting into the FSS error point just after an RF pi filter, but before the VGA. We wondered what the weird filter impedance was doing to our signal if we inject after it. I used LISO to model this FSS common section and attach the plots.

The first plot shows the TF between the Test 1 input and the AD602 VGA input. This is NOT the input that we are actually using.

The second plot shows the TF between the IN1 port (which we are actually using) and the VGA input.

Neither of them shows the 1 MHz bump that we see in the measurements, so I suspect that the board has been modified...the hunt continues. We've got to pop the top of the TTFSS and take photos and measure from IN1 to VGA input.

** FSScomm.fil is now in the LISO SVN. The following command line will run it with two different cases and cat the PDF files into one. If you use an auto-refresh PDF viewer like Okular or Mac Preview, its a nicer display than the usual GNUplot window:

./mfil FSScomm.fil; sleep 1; pdftk FSScomm_run*.pdf cat output FSScomm.pdf

Attachment 1: FSScomm.pdf
FSScomm.pdf FSScomm.pdf
  10379   Wed Aug 13 22:01:57 2014 ranaUpdateIMCCalculation for the input mode cleaner

Nic, Andres, and I discussed some more about the MC WFS project today. We want to shorten the proposed WFS2 path. Andres is going to explore moving the 2" diameter lens in coming up with layouts. We also want the WFS to face west so that we can see the diode face with an IR viewer easily and dump the reflected beams in the razor dumps.

We wondered about fixing the power levels and optical gain:

  1. What is the MC modulations depth? What would happen if we increase it a little? Does anyone know how to set it? Will this help the MC frequency noise?
  2. What is the max power on the WFS? I guess it should be set so that the power dissipation of the detector is less than 1 W with the MC unlocked. So P_diss = (100 V)*(I_tot), means that we should have less than 10 mA or ~50 mW when the MC is unlocked.
  3. Another consideration is saturation. The RF signals are tiny, but maybe the DC will saturate if we use any more power. The quadrants are saturated when unlocked and ~200 mV locked. According to D990249, the DC gain in the head is 1000 V/A. The measured power levels going into the heads (w/ MC unlocked) are: P_WFS1 = 4.9 mW and P_WFS2 = 7.7 mW. We don't have control of the DC gain, but there is a 10x and 100x switch available inside the demod board (D980233). From these numbers, I figure that we're in the 100x position and so the effective DC gain between photocurrent and the DC readback voltages is 100 kOhm. Therefore, we are in no danger of optical or electronics saturation. And the unlocked photocurrent of ~40/100000=0.4 mA => 0.04 W heat generated in the diode, so we're OK to increase the power level by another factor of 2-4 if we want.
  4.  We noticed that the ADC inputs are moving by ~50 counts out of 65000, so we're doing a really bad job of signal conditioning. This was previously noticed 6 years ago but we failed to follow up on it. Feh.

While checking this out, I converted the McWFS DC offsets script from csh to bash and committed it to the SVN. We need to remove the prefix 'feature' that Jamie has introduced to cdsutils so that we can use C1 again.

 

  10378   Wed Aug 13 19:23:09 2014 JenneUpdateLSCPSL, Aux laser mode hop check

This afternoon Q helped me put in some temporary PDs for checking for any mode hopping behavior in our 3 main lasers. 

Q helped me install PDA55s on each of the lasers (I did the ends, he did the PSL) so that we could do the mode hop temperature check.  For the Yend, I took the leakage transmission through the first Y1 steering mirror after the laser. This beam was dumped, so I replaced the dump with a PDA55. For the Xend, the equivalent mirrors are too close to the edge of the table, so I put in a spare Y1, and reflect most of the light to a beam dump.  The leakage transmission then goes to a PDA55.  Note that for both of these cases, no alignment of main laser path mirrors was touched, so we should just be able to remove them when we're through.  For the PSL, I believe that Q took the rejected light from one of the PBSes before the PMC. 

The end temporary PDs are using the TRX / TRY cables, so we will be looking at the C1:LSC-TR[x,y] channels for the power of the end lasers.  The PSL's temporary PD is connected to the PMC REFL cable.  For the end PDs, since I had filter banks available, I shuttered the end lasers and removed the dark offset.  I then changed the gains to 1, so the values are in raw counts.  The usual transmission normalization gains are noted in one of the control room notebooks.

I did a slow ezcastep and ramped the temperature of all 3 lasers over about an hour.  Since we usually use the PSL around FSS slow slider value of zero, I swept that from -10 to +10.  Since we usually use the Xend laser at around 10,000 counts, I swept that from 0 to 20,000.  For the Yend laser, it is usually around -10,000 counts, so I swept it from -20,000 to 0.  ezcastep -s 0.2 C1:ALS-X_SLOW_SERVO2_OFFSET +1,20000 C1:ALS-Y_SLOW_SERVO2_OFFSET +1,20000 C1:PSL-FSS_SLOWDC +0.001,20000

I was looking for something kind of similar to what Koji saw when he did this kind of sweep for the old MOPA (elog #2008), but didn't see any power jumps that looked suspicious.

Here is the PSL:

ModeHopCheck_PSL_13Aug2014.pdf

The Xend:

ModeHopCheck_Xend_13Aug2014.pdf

And the Yend:

ModeHopCheck_Yend_13Aug2014.pdf

  10377   Wed Aug 13 17:37:43 2014 JenneUpdateGeneralGame plan

2014_Aug_13.pdf

Here's the game plan for things that we need to do to get this IFO locked up. 

Red is for things that should be done today, or tomorrow if they don't get finished today (eg. laser mode hopping temperature check).  Orange is for things that will become red once the current red things are gone (eg. inferring the POP QPD gouy phase, and moving it to minimized PRM information).  Green is for things that we'd like to do, but aren't high priority (eg. X green mode matching).  Blue is for things that we should remember, but not plan on working on soon (eg. putting PZTs on the Yend table for green).

TODAY so far:

Q already did the tweak up of the PSL SHG crystal alignment.  HE SHOULD ELOG ABOUT THIS.  What was the final power of green that you got?  Do we have any record of a previous measurement to compare to?

Q helped me install PDA55s on each of the lasers (I did the ends, he did the PSL) so that we could do the mode hop temperature check.  For the Yend, I took the leakage transmission through the first Y1 steering mirror after the laser. This beam was dumped, so I replaced the dump with a PDA55. For the Xend, the equivalent mirrors are too close to the edge of the table, so I put in a spare Y1, and reflect most of the light to a beam dump.  The leakage transmission then goes to a PDA55.  Note that for both of these cases, no alignment of main laser path mirrors was touched, so we should just be able to remove them when we're through.  For the PSL, I believe that Q took the rejected light from one of the PBSes before the PMC.  He mentioned that he bumped something, so had to realign the beam into the PMC, but that he was able to get the transmission back up to 0.802, when we were seeing it in the mid 0.7's for the last several days.

The end temporary PDs are using the TRX / TRY cables, so we will be looking at the C1:LSC-TR[x,y] channels for the power of the end lasers.  The PSL's temporary PD is connected to the PMC REFL cable.  For the end PDs, since I had filter banks available, I shuttered the end lasers and removed the dark offset.  I then changed the gains to 1, so the values are in raw counts.  The usual transmission normalization gains are noted in one of the control room notebooks.

I did a slow ezcastep and ramped the temperature of all 3 lasers over about an hour.  I'll write a separate elog about how that went.

  10376   Wed Aug 13 16:12:55 2014 HarryUpdateGeneralFOL Layout Diagram

Per Q's request, I've made up a diagram of the complete FOL layout for general reference.

FOLLayout2.png

  10375   Wed Aug 13 13:08:24 2014 ranaUpdateIMCCalculation for the input mode cleaner

Can you please give us some more details on how this design was decided upon? What were the design considerations?

It would be nice to have a shorter path length for WFS2. What is the desired spot size on the WFS? How sensitive are they going to be to IMC input alignment? Are we still going to be recentering the WFS all the time?

  10374   Wed Aug 13 10:50:04 2014 AndresUpdateIMCCalculation for the input mode cleaner

  Calculation for the input mode cleaner

I have been working on the calculation for the input mode cleaner. I have come out with a new optical setup that will allow us increase the Gouy phase different between the WFS to 90 degrees. I use a la mode to calculate it. The a la mode solution :

   label            z (m)      type             parameters         
    -----            -----      ----             ----------         
    MC1                    0    flat mirror      none:            
    MC3               0.1753    flat mirror      none:            
    MC2              13.4587    curved mirror    ROC: 17.8700       
    Lens1            29.6300    lens             focalLength: 1.7183
    BS2              29.9475    flat mirror      none:            
    First Mirror     30.0237    flat mirror      none:            
    WFS1             30.2269    flat mirror      none:            
    Second Mirror    30.2650    flat mirror      none:            
    Third Mirror     30.5698    flat mirror      none:            
    Lens2            30.9885    lens             focalLength: 1     
    Fourth Mirror    31.0778    flat mirror      none:            
    Lens3            31.4604    lens             focalLength: 0.1000
    Fifth Mirror     31.5350    flat mirror      none:            
    Sixth Mirror     31.9414    flat mirror      none:            
    WFS2             31.9922    flat mirror      none:    
  

I attached a pictures how the new setup is supposed to look like. 

Attachment 1: ModeCleanerSetup0.PNG
ModeCleanerSetup0.PNG
Attachment 2: alaModeModeCleanersolution.png
alaModeModeCleanersolution.png
  10373   Wed Aug 13 10:49:39 2014 HarryUpdateGeneralWeekly Update

 In the past week, I designed and assembled coupling telescopes for the PSL and Y Arm Lasers

The Y Arm was coupled to ~5mV, and the PSL remains uncoupled.

 

For the next week, I'm planning on working on things like my presentation and/or final report.

Though as of last night, my computer refuses to turn on, so there may be some further "troubleshooting" involved in that whole process.

  10372   Wed Aug 13 03:03:37 2014 ericqUpdateGeneralGreen beatnote troubles

[Jenne, Rana, ericq]

No luck locking tonight, as spent a while trying to figure out the complete absence of the green beatnotes. Long story short, we ended up having to adjust the pointing on the PSL table.

Unrelated to this, we also turned on the noise eater on the PSL laser because why not. 


We hooked the BBPDs directly up to a 300MHz scope to try to see the beat as it happened. We witnessed a very strange intermittent ~800MHz oscillation on the Y BBPD, and weirder still, on both the RF and DC outputs of the PD, and the frequency was independent of the laser temperatures. This is to be investigated in the future, but was not related to the beat note state. 

Some progress was made when we took some components out, and looked at the far field of the PSL-Ygreen overlap, and saw some misalignment, and corrected it. Putting the end laser temperature in the usual area allowed the beat note to be found, with the eventual amplitude of ~-40dBm directly out of the BBPD. The Y green alignment was pretty bad throughout, so this can be improved to bring the beat amplitude up. We should also check and make sure we're well aligned to the SHG with the PSL light. We're leaving the X beat for tomorrow, now knowing that we should be able to get it with careful alignment. 

  10371   Tue Aug 12 23:07:24 2014 HarryUpdateGeneralPSL Telescope

I put the PSL telescope in place, and started coupling to it.

Unfortunately, I was only able to couple about 55 uW into the "fiber coupler" (read: fiber coupled splitter). See picture below:

PSLTelescopePic.png

Additionally, I'm not sure why this is, but both of the splitters we ordered don't split equally, but to 90% and 10% in each output port.

We also found that, since we aren't using the fibers we originally intended to, the specs are a little different, and the waist we're trying to have at the collimator face is now 283 um.

  10370   Tue Aug 12 18:20:13 2014 ericqUpdateIOOFSS box TFs

I made some measurements of the FSS box today, to have TFs for a loop model, but also to see what the difference between the different inputs was. 

As a reminder, the FSS box takes the error signal from the MC servo, does some filtering, and sends out two outputs: one to the laser PZT via KojiBox and Thorlabs HV amplifier, and one to be summed with the PMC modulation signal to the PC. Rana found the schematic at D040105

The MC error signal currently enters via a port called "IN1", but there is also a "Test 1 in," which experiences different filtering. I measured the TFs from each of these inputs to both the FAST and PC outputs. There is also an IN2, that is added after the offset point, but was not able to make a good measurement, for reasons unknown. From these TFs, I inferred the difference between the PC and FAST path, as well as the difference between IN1 and Test 1 in.

Specifically, I plugged the cable that is usually connected to the MC servo output, labelled "TO FSS BOX", into the RF out of the AG4395. I then took a BNC cable from the FAST out, or PC out, and fed it into a mini circuits DC block (BLK-89-S+), and then into input A, after checking on a scope that the signal was roughly zeroed and not too huge. Unbeknownst to me at the time, the PC drive output can be pretty big, and could potentially fry the analyzer's input. Fortunately, I think I avoided this fate. 

FSSbox.pdfFSSfilt.pdf

A ~1.3 MHz bump can be seen here, which would conspire with the bump in the demod board I measured yesterday, to steal even more phase around 1MHz. Maybe we can modify the FSS box to help our gain peaking situation out. 

The data is attached.

RXA: Shazam!

Attachment 3: FSSdata.zip
  10369   Tue Aug 12 14:29:01 2014 ericqUpdateGeneralReasonable alignment restored

I'm leaving the interferometer with the PRM aligned, so that all optics (except SRM) are near the center of their oplev range. I'm curious as to what their variance will be over the next day; this can inform whether we need to improve the ETMY oplev's angular range or not. 

 Here's an 12 hour minute-trend of all of the oplevs. The worst offenders are ITMY pitch and yaw, and ITMX pitch. 

Additionally, ETMY's yaw range is +-30urad, and here we see it wandering by 10 urad in a half day. We probably need more range.  

OLtrend.png

  10368   Tue Aug 12 13:31:58 2014 JenneUpdatePEMSeismometer cables in place, ready for sensors

[TaraV, Jenne]

The short cable from the slab to the sensor has been assembled and installed for the Trillium slab at the corner station.  The corner still needs the sensor and the long cable, both of which are in use by the gyro experiment.

The STS-2 cable that was running to the Xend was pulled, and the new long Guralp cable that Den made was installed with help from Andres.  The Xend just needs the sensor itself, which is also in use in gyro-land.

So, once we get the 2 seismometers and the one cable back from Zach, we should have 3 sensors nicely on the slabs that Den and Steve designed.

  10367   Tue Aug 12 02:09:39 2014 ericqUpdateGeneralReasonable alignment restored

I took over the IFO, after Jenne's locking efforts, which included manual alignment, since the ASS was doing bad things. 

For whatever reason, the Yarm ASS TT gains needed to be flipped back to go in the right direction. I've restored the old BURT snap file, and the ASS seems to work for now.  

Furthermore, I added some FMs to the Yarm ASS to be able to ramp down gains, to be done as new offsets are ramped in, so that a smooth offset transition is possible. The new version of the script works reasonably, but could be smoother still... Once I iron this out, I'll do the same change to the Xarm, and update the buttons. 

In any case, I was able to run ASS on both arms; single arm lock maxed out at around 0.85, maybe because we're only getting 0.78 from the PMC and 16k from the MC? I then aligned and locked the PRM, then reentered the oplevs on all of the PRMI optics. Oddly, the ETMs were at single uRads on their oplevs.

With this arm alignment, I was able to get the green TRX to ~0.55, and thus the beatnote to around -25dBm, which is still lower than we'd like. I didn't touch the Y green alignment, though it is pretty bad, at transmission of below 0.2 when "locked" on the 00 mode. 

When I try to lock things, the initial ALS CARM and DARM locking seems to go fine, actuating on the ETMs for both DoFs, but ETMX is getting kicked during the resonance search every time. Maybe improving green alignment / increasing beatnote amplitudes will hopefully help some.

I'm leaving the interferometer with the PRM aligned, so that all optics (except SRM) are near the center of their oplev range. I'm curious as to what their variance will be over the next day; this can inform whether we need to improve the ETMY oplev's angular range or not. 

  10366   Mon Aug 11 23:50:38 2014 ranaConfigurationWikiDokuWikis are back up

Quote:

Quote:

It looks like auth is broken on the AIC wiki (though working fine on ATF and Cryo). I did some poking around but can't see how anything we did could have broken it.

I went into local.php and changed $conf['useacl'] = 1; to $conf['useacl'] = 0; and it looks like the auth issue goes away (I've changed it back). This isn't a fix (we want to use access control), but it gives us a clue as to where the problem is.

 There was still some residual permissions issue. This is now bypassed and so the ACL is ON and all seems to be back the way it was. I've tested that I can login and edit the wiki.


Some useless knowledge follows here. Please ignore.

After some hours of reading unhelpful DokuWiki blogs, I just put the backup wiki into the local disk on NODUS and then made a soft link to point to that from /users/public_html/wiki/. So this implies that the new NFS setup on chiara is different enough that it doesn't allow read/write access to the apache user on the NODUS/Solaris machine.

  10365   Mon Aug 11 23:32:54 2014 ericqSummaryIOOMC demod measurement

Here's the magnitude plot of the board TF. As mentioned above, this was done with Marconi+Scope, so we were not able to get the phase of this transfer function. 

MCDemod.pdf

Oddly enough, the bump that I saw is not included in Minicircuit's data on the SCLF-5.

Attachment 2: demodLP.txt
# F(Hz) RMS(mV)
1035 38.6
2031 38.47
4031 38.47
8032 38.38
16030 38.10
32030 38.10
64030 38.16
128000 38.10
256000 38.22
... 12 more lines ...
  10364   Mon Aug 11 22:07:31 2014 KojiSummaryIOOMC demod measurement

SCLF-5!? It's surprising as the cut off of the OLTF is just above 1Hz. cf this entry

This means that not the demod board but MC or FSS boards seem to have large attenuation above 1MHz.

In this situation, does SCLF-10/10.7 really help us?

  10363   Mon Aug 11 21:03:48 2014 ericq, ranaSummaryIOOMC demod measurement

We measured the TF of the MC Demod board today.

We set the Marconi to +3dBm and drove the PD IN port of the demod board, starting at 29.5 MHz. Then we looked at the beat signal amplitude in the output of the demod board. So this is a transfer function but with mag only. Plots from Q below.

Rana took the demod board out and took pictures of it. Inside, the post mixer low pass is a SCLF-5 from mini-circuits. This has a lot of cutoff down low. Since the purpose of this filter is only to cutoff the 2f-1f and the 3f-2f products, we need to have a lot of attenuation at 29.5 MHz. One day, we may want to re-instate that notch for the (3*f1- f_MC) beat frequency, but for now we want stability.

So, I recommend that we (Steve) get 3 each of the SCLF-10 and SCLF-10.7 from Mini-Circuits Tuesday morning. Maybe we can put them into a spare board?

Also, we should probably remove the 140kHz:70kHz lead filter which is in the MC servo board. Its out of date. I think it would be fine for us to get a 7-15 kHz UGF for the CM servo and the MC can basically do that already. Mainly we want to fix the high frequency shape to get more stability.

After the measurements and photos, we had to reset the MCWFS offsets to get the WFS to not break the lock. Seems very sensitive to offsets. Hopefully Andres will give us a new Gouy phase telescope.

  10362   Mon Aug 11 10:23:39 2014 steveSummarySUSoplev laser summary updated

Quote:

 

                  2005              ALL oplev servos use Coherent DIODE LASERS # 31-0425-000, 670 nm, 1 mW

    Sep. 28, 2006              optical lever noise budget with DC readout in 40m,  LIGO- T060234-00-R, Reinecke & Rana

    May  22, 2007              BS, SRM & PRM  He Ne 1103P takes over from diode

    May  29, 2007              low RIN He Ne JDSU 1103P selected, 5 purchased sn: T8078254, T8078256, T8078257, T8078258 & T8077178 in Sep. 2007

    Nov  30, 2007               Uniphase 1103P divergence measured

    Nov. 30, 2007               ETMX old Uniphase 1103P  from 2002 dies: .............., running time not known......~3-5 years?

    May 19, 2008               ETMY old Uniphase 1103P from 1999 dies;.....................running time not known.....~    ?

    Oct.  2, 2008                ITMX & ITMY are still diodes, meaning others are converted to 1103P earlier

 

                     JDSU 1103P were replaced as follows:

   May 11, 2011                ETMX replaced, life time 1,258 days  or 3.4 years

   May 13, 2014               ETMX , LT 1,098 days or 3 y

   May 22, 2012               ETMY,  LT 1,464 days or  4 y

   Oct.  5, 2011                BS & PRM, LT 4 years,  laser in place at 1,037 days or 2.8 y

   Sep. 13, 2011               ITMY  old 1103P &    SRM    diode laser replaced by 1125P  ..........old He life time is not known, 1125P in place 1,059 days or 2.9 y

   June 26, 2013              ITMX 622 days or 1.7 y    note: we changed because of beam quality.........................laser in place 420 days or 1.2 y

 

  Sep. 27, 2013               purchased 3 JDSU 1103P lasers, sn: P893516, P893518, P893519 ......2 spares ( also 2 spares of 1125P of 5 mW & larger body )

 

      May  13, 2014             ETMX,  .............laser in place 90 d

      May  22, 2012             ETMY, 

     Oct.  7,  2013             ETMY,  LT  503 d  or  1.4 y............bad beam quality ?

     Aug. 8,  2014              ETMY,  .............laser in place   425 days  or  1.2 y

 

  10361   Mon Aug 11 08:04:01 2014 SteveUpdateSUSETMX damping restored
  10360   Sun Aug 10 00:54:54 2014 HarryUpdateGeneralAUX Couping

The Y End laser dumped SHG light has been coupled into the yellow fiber that terminates at the PSL table.

It's not super stably coupled, and only at 5mW. I'll be interested to see what it is on monday.

  10359   Sat Aug 9 14:35:28 2014 KojiSummaryIOOMC servo analysis

Eric's OLTF turned out consistent with the AO path TF that has been measured by me on Jul 31 (entry 10322).

Attachment 1:
Updated empirical fit of the open loop TF by LISO.
In this fit, I gave some of the poles/zeros associated with the boost manually set so that I can use them for the servo design.
LISO itself can make better fitting if all of the variables are moved.

Atatchment 2:
The OLTF data and LISO source for the fitting.

Attachment 3:
Comparison of the AO path TFs. The red one was measured directly on Jul 31. The TF is normalized at the low frequency.
The blue was estimated from the OLTF model given above. They are well consistent now.

Attachment 4:
Now some servo design was tried. In the new design (blue), zeros of the super boost frequency was moved from 20kHz to 30kHz
with the hope of having flatter AO response. The improvement is very little while costing costing above 100kHz. Note that the vertical
axis is intentionally in a linear scale. In fact, the AO response is much improved compared to the one before the MC UGF was increased
(shown in magenta). We have a flatter response both in magnitude and phase.
Therefore I think there is no need to tweak the boost frequency for the AO path.
I'd rather recommend to inspect the high frequency LPFs to earn more gain margin at 1MHz as
explained in entry 10322.

Attachment 5:
This figure shows the comparison of the TFs for the current and new design trial, just in case someone is interested in to see.

 

Attachment 1: MC_OLTF_Fit.pdf
MC_OLTF_Fit.pdf
Attachment 2: liso.zip
Attachment 3: MC_CLTF_Fit.pdf
MC_CLTF_Fit.pdf
Attachment 4: MC_CLTF_new.pdf
MC_CLTF_new.pdf
Attachment 5: MC_OLTF_new.pdf
MC_OLTF_new.pdf
  10358   Fri Aug 8 20:22:12 2014 JenneMetaphysicsGeneralkitchen sink water off

Quote:

 When I got back to the lab, there was enough water that it was seeping under the wall, and visible outside. Physical plant says it will take an hour before they can come, so I'm getting dinner, then will let them in.

 The guy from physical plant came, and turned off the water to the kitchen sink.  He is putting in a work order to have the plumbers come look at it on Monday morning.  It looks like something is wrong with the water heater, and we're getting water out of the safety overpressure valve / pipe.

The wet things from under the sink are stacked (a little haphazardly) next to the cupboards.

  10357   Fri Aug 8 19:42:59 2014 JenneMetaphysicsGeneralkitchen sink flooding

 When I got back to the lab, there was enough water that it was seeping under the wall, and visible outside. Physical plant says it will take an hour before they can come, so I'm getting dinner, then will let them in.

  10356   Fri Aug 8 18:08:12 2014 KojiSummaryIOOMC servo analysis

The closed gain I meant is the AO path: Use IN2 to excite the MC loop and measure IN1 using MON2(?).
In order to obtain the open loop gain from this meausrement, the gain mismatching needs to be compensated, though.

This measurement is to correctly predict the AO path response from the open loop transfer function.

Anyway, the openloop gain seems nicely measured. I'll try to predict AO path response from this.

  10355   Fri Aug 8 16:45:40 2014 NichinUpdateWikiPDFR wiki updated

 The PDFR system has been documented in the 40m wiki and all the relevant information about making changes and keeping it updated have been mentioned.

https://wiki-40m.ligo.caltech.edu/Electronics/PDFR_system

This pretty much wraps up my SURF 2014 project at the 40m lab. 

  10354   Fri Aug 8 15:57:29 2014 ericqSummaryIOOMC servo analysis

 I did some further measurements, to try and see what corresponds to what. In the end I performed four measurements:

  1. Closed loop gain measurement on SR785: Source to MC exc, T'd to channel one. Test 2 to channel two.
  2. Open loop gain measurement on SR785: Source to MCexc, Test 2 to channel one, Test 1 to channel two.
  3. Closed loop gain measurement on AG4395: RF Source to MC exc, T'd to R input. Test 2 to A input.
  4. Open loop gain measurement on AG4395: RF Source to MC exc, Test 2 to R input. Test 1 to A input.

I then converted OLGs to CLG and vice-versa with CLG = 1/(1-OLG)

Here are two plots showing the measured and inferred loop TFs for both closed and open. 

OLTFs.pdfCLTFs.pdf

The best agreement seems to be between the directly measured OLGs. Maybe I did something weird with the CLG measurements, or input impedances are distorting things ... 

All data is attached, along with code used to generate the plots. 

Attachment 3: mcLoopAug8.zip
  10353   Fri Aug 8 14:42:41 2014 AkhilUpdateGeneralPID loop Design for beat note stabilization

 The attached in a zip file are the Simulink feedback loop models for the FOL for both X and Y ends. The controller PID values are estimated by setting a temperature count reference point to 5344, which corresponds to 100 MHz frequency.  The plant transfer function is as calculated in my previous elogs.

 We were not  able to test the PID loop , with the green laser by PZT actuation because of the misalignment of the arms and non-existence of the beat note since last few days. However, we have a complete idea of the design and PID parameters that will be used for the FOL with infrared laser. So we decided that it would be better to test the loop by temperature actuation after the fiber optics is installed and the coupling of infrared laser into the fiber is complete. As of now, we have planned to place the FOL box inside so that it can be used to obtain the green laser beat note on the StripTool graphs. 

Attachment 1: PID.zip
  10352   Fri Aug 8 14:27:18 2014 AkhilUpdateComputer Scripts / ProgramsFOL Scripts

 The scripts written for interfacing the FC with R Pi, building EPICS database, piping data into EPICS channels,PID loop for FOL are contained in :

 /opt/rtcds/caltech/c1/scripts/FOL 

The instructions to run these codes on R Pi( controls@domenica) will be available on FOL 40m wiki page.

Also instructions regarding EPICS installation on R Pi and building an EPICS SoftIoc to streamline data from hardware devices into channels will be updated shortly.

 

 

 

 

 

 

  10351   Fri Aug 8 12:39:19 2014 ericqSummaryIOOMC servo analysis

I have measured the current boosted MC CLG below 100kHz with an SR785. Swept sine only could get me down to 10kHz, but I was able to get down to 5kHz with a noise-injection measurement. 

MCloopAug8.pdf

I am attaching the SR785 outputs, which are in dB and Degrees. Additionally I pruned the areas of bad coherence out of these, and merged them to provide data files for the CLG and OLG in Real,Imaginary format.

Attachment 1: mcLoopAug8.zip
  10350   Fri Aug 8 11:22:35 2014 steveSummarySUSoplev laser summary

 

                  2005              ALL oplev servos use Coherent DIODE LASERS # 31-0425-000, 670 nm, 1 mW

    Sep. 28, 2006              optical lever noise budget with DC readout in 40m,  LIGO- T060234-00-R, Reinecke & Rana

    May  22, 2007              BS, SRM & PRM  He Ne 1103P takes over from diode

    May  29, 2007              low RIN He Ne JDSU 1103P selected, 5 purchased sn: T8078254, T8078256, T8078257, T8078258 & T8077178 in Sep. 2007

    Nov  30, 2007               Uniphase 1103P divergence measured

    Nov. 30, 2007               ETMX old Uniphase 1103P  from 2002 dies: .............., running time not known......~3-5 years?

    May 19, 2008               ETMY old Uniphase 1103P from 1999 dies;.....................running time not known.....~    ?

    Oct.  2, 2008                ITMX & ITMY are still diodes, meaning others are converted to 1103P earlier

 

                     JDSU 1103P were replaced as follows:

   May 11, 2011                ETMX replaced, life time 1,258 days  or 3.4 years

   May 13, 2014               ETMX , LT 1,098 days or 3 y

   May 22, 2012               ETMY,  LT 1,464 days or  4 y

   Oct.  5, 2011                BS & PRM, LT 4 years,  laser in place at 1,037 days or 2.8 y

   Sep. 13, 2011               ITMY  old 1103P &    SRM    diode laser replaced by 1125P  ..........old He life time is not known, 1125P in place 1,059 days or 2.9 y

   June 26, 2013              ITMX 622 days or 1.7 y    note: we changed because of beam quality.........................laser in place 420 days or 1.2 y

 

  Sep. 27, 2013               purchased 3 JDSU 1103P lasers, sn: P893516, P893518, P893519 ......2 spares ( also 2 spares of 1125P of 5 mW & larger body )

 

  10349   Thu Aug 7 17:09:53 2014 HarryUpdateGeneralAUX Coupling In Progress

 I'm currently in the process of coupling dumped SHG light from the Y arm end table into fibers for FOL.

The main point is that the NPRO at that end in shuttered, because I wasn't sure whether or not leaving it open would've set anything on fire.

  10348   Thu Aug 7 16:47:35 2014 ericqUpdateSUSOplev Checkup

 I noticed some weird behavior on the ETMY oplev that led me to check them all out. 

The short of it is that the ETMY oplev has a pretty small angular range, compared to the displays and other oplevs. I measured how much angular motion each oplev can sense before the beam no longer hits all four quadrants (thus losing the ability to sense).  This could account for some of the additional angular motion of the mirrors... maybe. 

Also, some of the QPD quadrants had offsets as big as 400 counts, thus distorting the zero point. Anyways, here are the angular ranges of each QPD, assuming the current urad/cnt calibrations are valid. 

EMTY

  • P: +- 25urad
  • Y +- 30urad

ITMY

  • P:+-160urad
  • Y:+-172urad

 

BS

  • P:+-43urad
  • Y:+-40urad

 

ITMX

(Note: ITMX's oplev pitch and yaw is almost 30 degrees off of the alignment sliders' pitch/yaw coordinates. Steve tells me this is due to the tight nature of getting the oplev beam to the mirror without clipping.)

  • P:+-110urad
  • Y:+-80urad

 

ETMX

  • P:+-45urad
  • Y:+-85urad

 

PRM

  • P:+-50urad
  • Y:+-45urad

 

SRM

  • P:+-80urad
  • Y:+-80urad

I wrote a script to zero all of the QPD quadrants' offsets (it lives in /scripts/OL) and have used it successfully. The oplev laser must  be off before using it. 

  10347   Thu Aug 7 14:50:43 2014 HarryUpdateGeneralPER Measurement

 Purpose

I wanted to do a more robust measurement of PER of PM fibers for FOL, so I thought up this scheme.

Methods

I put together a setup as depicted below in order to take measurements of PER.

PERFinalSetup.png

The first thing to do was to calibrate the whole setup. In order to do so, I first used the quarter and half wave plates closest to the NPRO to eliminate as much ellipticity from the output beam as possible, and then rotate the newly linearized light to be in alignment with the transmittance of the first polarizing beam splitter (P-Polarization).

I then aligned the fiber's fast axis with the P-Polarization on both the input and output sides. This was important so that no virtual ellipticity would be measured in the final measurement of PER.

I then mode matched and fiber coupled the first PBS output into the fibers, to about 30 mW (~60% coupling).

Photodiode Calibration

I wanted to measure both intensity of P and S simultaneously, so as to minimize the random little time-varying changes that would affect the measurements, so I used a powermeter and a PD, calibrated with the aformentioned powermeter.

In order to be able to compare the photodiode (PDA520) output to the powermeter (Orion) output, I fixed them each in their positions, and varied the laser power to produce the type of linear relationship we expect to see between PD Voltage and Optical Power. In this case, the conversion was P = V*2.719.

PDCalibration.png

PER Measurement

As opposed to the first method, which took only one datum, this method records P and S simultaneously, at different points through rotation of a linearly polarized beam.

Using the second HWP, I rotated the linearly polarized beam before it entered the fiber, at each point, recording the outputs of the PD and the Powermeter.

These data were then converted to be the same units, and fit to a sine wave.

Polarization_Intensity_Variation.png

As you can see, the intensities vary nearly identically, at a half wavelength phase difference, which is what one expects in this case. The PER of each polarization can be calculated by dividing the maximum value of one by the minimum of the other, and vice versa. The fact that these oscillate as we expect shows that the beam is relatively well linearized, and essentially that everything is working as it is assumed to be.

By looking at these fits, however, it is visible that they do not overlap with the actual extrema of the data. So, in order to produce more realistic values of extrema, those particular regions were fit to second order polynomials.

Extrema.png

The values of these extrema yield the following measurements:

(SMin / PMax) = 0.007 +/- .004  --->  -21.54 +/- 2.48 dB

(PMin / SMax) = 0.022 +/- .009  --->  -16.58 +/- 1.78 dB

Conclusion

The problem I find with these measurements is that they're hard to reproduce.

Plus they seem high, since non-PM fibers advertise extinction ratios around -30 dB., plus I measured it at roughly -24 dB the first time I tried.

Moving Forward

 

The next thing to do in terms of fiber characterization is to measure the frequency noise they introduce.

With respect to FOL, I just need some time to work on the PSL table, and at the Y end to couple the dumped SHG light, and then we can start using 1064nm beat notes to test//implement the feedback control system.

Attachment 5: PEReport.zip
  10346   Thu Aug 7 13:39:41 2014 NichinUpdateComputer Scripts / ProgramsWrapping up PDFR

1)The PDFR scripts have all been migrated into /scripts/PDFR/

2) The MEDM screen to run PDFR is /medm/MISC/PDFR.adl

3) A new button has been added on sitemap to open the above medm window.

4) All data and plots generated will sit in /scripts/PDFR/"PD Name"/

5) All features are working after the migration and absolute file paths are being used.

Work Remaining : Manual for others to make changes and keep using my system.

 

  10345   Thu Aug 7 12:34:56 2014 JenneUpdateLSCSuspensions not kicking?

Yesterday, Q helped me look at the DACs for some of the suspensions, since Gabriele pointed out that the DACs may have trouble with zero crossings.  

First, I looked at the oplevs of all the test masses with the oplev servos off, as well as the coil drive outputs from the suspension screen which should go straight out to the DACs.  I put some biases on the suspensions in either pitch or yaw so that one or two of the coil outputs was crossing zero regularly.  I didn't see any kicks. 

Next, we turned off the inputs of the coil driver filter banks, unplugged the cable from the coil driver board to the satellite box, and put in sinusoidal excitations to each of the coils using awggui.  We then looked with a 'scope at the monitor point of the coil driver boards, but didn't see any glitches or abnormalities.  (We then put everything back to normal)

Finally, I locked and aligned the 2 arms, and just left them sitting.  The oplev servos were engaged, but I didn't ever see any big kicks. 

I am suspicious that there was something funny going on with the computers and RFM over the weekend, when we were not getting RFM connections between the vertex and the end stations, and that somehow weird signals were also getting sent to some of the optics.  Q's nuclear reboot (all the front ends simultaneously) fixed the RFM situation, and I don't know that I've seen any kicks since then, although Eric thinks that he has, at least once.  Anyhow, I think they might be gone for now.

  10344   Thu Aug 7 12:25:14 2014 JenneUpdateIOOFSS offset changed

Quote:

The fast feedback should be around zero now!

 Dang it, I completely forgot.  Well, anyhow, it pulled itself back down to less than 1V, and the MC stayed happy for several hours.  I'm not totally sure what changing the offset did, but the MC seems happy for right now.  I should take a quick look at the error point to make sure that I didn't mess up your tuning.

  10343   Thu Aug 7 11:57:59 2014 KojiSummaryIOOMC servo analysis

LISO Fit for the IMC open loop TF. The data and liso source for the fitting were attached in the ZIP file.

I noticed now that the open loop TF I measured has too less phase delay.
I used the closed loop TF to estimate the openloop TF.

Looking at this comparison, I'm afraid that the superboost was not on during the measurement.
I need a new measurement to design MC loop modification to give the AO path for broader bandwidth.

Attachment 1: MC_OLTF_Fit.pdf
MC_OLTF_Fit.pdf
Attachment 2: IMC_OLTF.zip
Attachment 3: MC_OLTF_estimated.pdf
MC_OLTF_estimated.pdf
  10341   Wed Aug 6 21:22:09 2014 KojiUpdateIOOFSS offset changed

The fast feedback should be around zero now!

  10340   Wed Aug 6 17:29:36 2014 JenneUpdateIOOFSS offset changed

The MC has been unstable and unhappy for the last several hours.  When I looked, I saw that the FSS_FAST monitor has been hovering around 1 V, when it is supposed to be closer to 5ish. 

I changed the C1:PSL-FSS_INOFFSET from -0.08 to -0.8537, and will see if the MC sticks around for longer this time around.

  10339   Wed Aug 6 13:17:21 2014 ericqOmnistructureCDScdsutils: multifarious upgrades

I've checked out cdsutils-274 to /opt/rtcds/cdsutils, and updated the /ligo/apps/ligoapps-user-env.sh to have the newer machines use it by default. This was to gain access to the cdsutils.Step methods for use in the smooth ASS handoffs script. 

  10338   Wed Aug 6 12:44:52 2014 KojiUpdateASCPOP QPD signals

This is nice. Can we test this idea with POP22 + a razor blade?

Just to take transfer functions in PRMIsb between the PRM angle to POP QPD/POP22+razor blade
as well as the noise spectrum measurement are already useful.

We want to figure out the requirement for the 2f QPD.
(Transimpedance / Noise level / Beam size / etc)

Depending on the requirement we'll see if we need demodulation or just a power detector.

  10337   Wed Aug 6 10:45:56 2014 GabrieleUpdateASCPOP QPD signals

 In addition to the simulation described in my previous elog, I simulated the signal on a quadrant photodetector demodulated at 2F. The input laser beam is modulated at 11MHz up to the fifth order. There is no additional 55 MHz modulation.

The QPD demodulated at 2F shows good signals for PRC control for all CARM offsets, as expected from the previous simulation.

pop_qpd_2f.png

 

  10336   Wed Aug 6 10:10:45 2014 HarryUpdateGeneralWeekly Plan 8.6.14

Last Week

 

Took first round of PER measurements after a long setup.

Started setting up to take measurement of the other polarization--ran into issues with mounts again. (Spinning of their own free will again.)

Devised a new scheme for taking more robust measurements of PER--still in progress.

Next Week

Finish data analysis of these latest PER measurements

Hopefully finally move on to frequency noise characterization

Materials Needed

None for PER

Unknown for frequency noise

 

  10335   Wed Aug 6 00:14:10 2014 JenneUpdateLSCALS is iffy tonight

The ALS system is iffy tonight.

After putting the cable back to the RF spectrum analyzer (it had been taken to test the frequency counter setup, and not put back), I had a good Yarm beatnote, but again this evening the Xarm beatnote is small.  I touched up the PSL table alignment (very, very little needed, but it did double my peak height).  I *think* that this is happening because we haven't settled into a good IFO alignment place, so the arm pointing keeps changing very slightly, which means that the PSL ALS alignment needs touching.  Anyhow, even after alignment the Xarm beatnote is only -36 dBm at 81 MHz.  It should be at least -25 dBm or so, although I haven't seen it any larger than about -35 dBm since the IFO beam was lost last Friday.

I am not able to hold ALS lock long enough to scan the arms and find the IR resonances.  The only optics that I am actuating on this evening are the 2 ETMs.  When I lose lock and look at the watchdogs, the ETMs are the only optics that have largeish numbers, which comes from the ALS lockloss.  So, I don't think I am suffering from the ITM suspension kicks tonight.  Rather, I think that it's that the ALS system isn't tuned up nicely.

I think that it is past time we tuned up and checked out the ALS PDH setup.  Q:  Can you please measure the loop TFs for both of the ALS PDH boxes tomorrow?  At the very least we want to know what we're working with. 

Evan:  What is the status with the ISS? 

I am going to try tomorrow to look at the suspensions, and see if I can track anything down.  I feel like I see the kicks more often when the arms are locked, i.e. we are sending an LSC signal to them.  The LSC POS signal is a factor of a few hundred larger than the damping SUSPOS signal is.  Are we saturating something somewhere?  Why is this a new thing?  We certainly do see kicks when the LSC is not engaged, so this may not be the right path, but it is something concrete to look at.

  10334   Tue Aug 5 19:20:05 2014 AkhilUpdateGeneralPID loop Design for beat note stabilization

 Today I and EricQ went inside the lab and set up the cables running from the a DAC channel into  PZT input so that we can use the PID controller to tune in the PZT offset to maintain the beat note within a detectable range (This is plan B as the main plan of actuating on the laser temperature can be achieved only after the fiber setup with the PSL is ready). I obtained all the poles and zeroes of plant and started designing a PID loop to test it with the existing system.

I will put in my PID values into the already existing PERL controller code (that is used for controller design in the 40m) and run tests with the PID loop while actuating on the PZT offset. 

 

ELOG V3.1.3-